a charge of 5c and 3kg mass is in vertical equilibrium due to electric field E find the electric field E.
1
Expert's answer
2020-09-29T09:42:37-0400

According to the second Newton's law, gravitational force and Lorenz force must compensate each other.

"mg = qE"

"\displaystyle E=\frac{mg}{q}= \frac{3 \cdot 9.81}{5}=5.886 \approx 5.9 \, \frac{V}{m}"

Answer: "5.9 \, \frac{V}{m}"